(a) Is an aqueous solution of LiHSO4 acidic, basic, or neutral? (b) Write the net ionic...

(a)

Is an aqueous solution of LiHSO4 acidic, basic, or neutral?

(b)

Write the net ionic reaction that occurs between LiHSO4 and water. (Omit states-of-matter from your answer.)

(c)

Calculate the pH of a 0.109 M solution of LiHSO4.
Answer #1

(a) Acidic

Explanation: LiHSO4 still has one replaceable hydrogen.

(b) LiHSO4(aq) + H2O(l) ----> LiOH(aq) + H2SO4(aq)

(c) pKa of HSO4- = 1.921

Now, pH = 1/2 (pKa - Log[NaHSO4]

i.e. pH = 1/2 {1.921 - Log(0.109)}

i.e. pH = 1/2 {1.921 - (-0.963)}

Therefore, pH = 1.442